In an examination 7% of students score less than 35% marks and 89% of students score less than 60% marks. Find the mean and standard deviation, if the marks are normally distributed.


1
Expert's answer
2021-08-03T15:18:40-0400

Let "X=" student's mark: "X\sim N(\mu, \sigma)"


"P(X<x_1)=P(Z<\dfrac{x_1-\mu}{\sigma})=0.07"

"P(X<x_2)=P(Z<\dfrac{x_2-\mu}{\sigma})=0.89"

"\dfrac{x_1-\mu}{\sigma}\approx-1.475791"

"\dfrac{x_2-\mu}{\sigma}\approx1.226528"

"x_1=0.35, \dfrac{0.35-\mu}{\sigma}\approx-1.475791"

"x_2=0.6, \dfrac{0.6-\mu}{\sigma}\approx1.226528"

"\mu=0.35+1.475791\sigma"

"\mu=0.6-1.226528\sigma"

"0.35+1.475791\sigma=0.6-1.226528\sigma"

"\sigma=\dfrac{0.6-0.35}{1.475791+1.226528}\approx0.0925"

"\mu\approx0.4865"

"\mu=48.65\%, \sigma=9.25\%"